\begin{thebibliography}{51}
\providecommand{\natexlab}[1]{#1}

\bibitem[{AI@Meta(2024)}]{llama3modelcard}
AI@Meta. 2024.
\newblock \href {https://github.com/meta-llama/llama3/blob/main/MODEL_CARD.md}
  {Llama 3 model card}.

\bibitem[{Ali et~al.(2024)Ali, Fromm, Thellmann, Rutmann, L{\"u}bbering,
  Leveling, Klug, Ebert, Doll, Buschhoff, Jain, Weber, Jurkschat, Abdelwahab,
  John, Ortiz~Suarez, Ostendorff, Weinbach, Sifa, Kesselheim, and
  Flores-Herr}]{ali-etal-2024-tokenizer}
Mehdi Ali, Michael Fromm, Klaudia Thellmann, Richard Rutmann, Max
  L{\"u}bbering, Johannes Leveling, Katrin Klug, Jan Ebert, Niclas Doll, Jasper
  Buschhoff, Charvi Jain, Alexander Weber, Lena Jurkschat, Hammam Abdelwahab,
  Chelsea John, Pedro Ortiz~Suarez, Malte Ostendorff, Samuel Weinbach, Rafet
  Sifa, Stefan Kesselheim, and Nicolas Flores-Herr. 2024.
\newblock Tokenizer choice for {LLM} training: Negligible or crucial?
\newblock In \emph{Findings of the Association for Computational Linguistics:
  NAACL 2024}, pages 3907--3924.

\bibitem[{Almazrouei et~al.(2023)Almazrouei, Alobeidli, Alshamsi, Cappelli,
  Cojocaru, Debbah, Étienne Goffinet, Hesslow, Launay, Malartic, Mazzotta,
  Noune, Pannier, and Penedo}]{almazrouei2023falcon}
Ebtesam Almazrouei, Hamza Alobeidli, Abdulaziz Alshamsi, Alessandro Cappelli,
  Ruxandra Cojocaru, Mérouane Debbah, Étienne Goffinet, Daniel Hesslow,
  Julien Launay, Quentin Malartic, Daniele Mazzotta, Badreddine Noune, Baptiste
  Pannier, and Guilherme Penedo. 2023.
\newblock \href {https://arxiv.org/abs/2311.16867} {The falcon series of open
  language models}.

\bibitem[{Bisk et~al.(2020)Bisk, Zellers, Le~bras, Gao, and
  Choi}]{Bisk_Zellers_Lebras_Gao_Choi_2020}
Yonatan Bisk, Rowan Zellers, Ronan Le~bras, Jianfeng Gao, and Yejin Choi. 2020.
\newblock Piqa: Reasoning about physical commonsense in natural language.
\newblock pages 7432--7439.

\bibitem[{Black et~al.(2022)Black, Biderman, Hallahan, Anthony, Gao, Golding,
  He, Leahy, McDonell, Phang, Pieler, Prashanth, Purohit, Reynolds, Tow, Wang,
  and Weinbach}]{black2022gptneox20b}
Sid Black, Stella Biderman, Eric Hallahan, Quentin Anthony, Leo Gao, Laurence
  Golding, Horace He, Connor Leahy, Kyle McDonell, Jason Phang, Michael Pieler,
  USVSN~Sai Prashanth, Shivanshu Purohit, Laria Reynolds, Jonathan Tow, Ben
  Wang, and Samuel Weinbach. 2022.
\newblock \href {https://arxiv.org/abs/2204.06745} {Gpt-neox-20b: An
  open-source autoregressive language model}.

\bibitem[{Brown et~al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal,
  Neelakantan, Shyam, Sastry, Askell, Agarwal, Herbert-Voss, Krueger, Henighan,
  Child, Ramesh, Ziegler, Wu, Winter, Hesse, Chen, Sigler, Litwin, Gray, Chess,
  Clark, Berner, McCandlish, Radford, Sutskever, and
  Amodei}]{NEURIPS2020_1457c0d6}
Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared~D Kaplan, Prafulla
  Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell,
  Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon
  Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, Chris
  Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess,
  Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever,
  and Dario Amodei. 2020.
\newblock Language models are few-shot learners.
\newblock In \emph{Advances in Neural Information Processing Systems 33
  (NeurIPS)}, pages 1877--1901.

\bibitem[{Chelba et~al.(2013)Chelba, Mikolov, Schuster, Ge, Brants, Koehn, and
  Robinson}]{41880}
Ciprian Chelba, Tomas Mikolov, Mike Schuster, Qi~Ge, Thorsten Brants, Phillipp
  Koehn, and Tony Robinson. 2013.
\newblock \href {http://arxiv.org/abs/1312.3005} {One billion word benchmark
  for measuring progress in statistical language modeling}.
\newblock Technical report, Google.

\bibitem[{Chen et~al.(2019)Chen, Su, Shen, Chen, Yan, and
  Wang}]{chen-etal-2019-large}
Wenhu Chen, Yu~Su, Yilin Shen, Zhiyu Chen, Xifeng Yan, and William~Yang Wang.
  2019.
\newblock How large a vocabulary does text classification need? a variational
  approach to vocabulary selection.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies (NAACL-HLT)}, pages 3487--3497.

\bibitem[{Chowdhery et~al.(2022)Chowdhery, Narang, Devlin, Bosma, Mishra,
  Roberts, Barham, Chung, Sutton, Gehrmann, Schuh, Shi, Tsvyashchenko, Maynez,
  Rao, Barnes, Tay, Shazeer, Prabhakaran, Reif, Du, Hutchinson, Pope, Bradbury,
  Austin, Isard, Gur-Ari, Yin, Duke, Levskaya, Ghemawat, Dev, Michalewski,
  Garcia, Misra, Robinson, Fedus, Zhou, Ippolito, Luan, Lim, Zoph, Spiridonov,
  Sepassi, Dohan, Agrawal, Omernick, Dai, Pillai, Pellat, Lewkowycz, Moreira,
  Child, Polozov, Lee, Zhou, Wang, Saeta, Diaz, Firat, Catasta, Wei,
  Meier-Hellstern, Eck, Dean, Petrov, and Fiedel}]{chowdhery2022palm}
Aakanksha Chowdhery, Sharan Narang, Jacob Devlin, Maarten Bosma, Gaurav Mishra,
  Adam Roberts, Paul Barham, Hyung~Won Chung, Charles Sutton, Sebastian
  Gehrmann, Parker Schuh, Kensen Shi, Sasha Tsvyashchenko, Joshua Maynez,
  Abhishek Rao, Parker Barnes, Yi~Tay, Noam Shazeer, Vinodkumar Prabhakaran,
  Emily Reif, Nan Du, Ben Hutchinson, Reiner Pope, James Bradbury, Jacob
  Austin, Michael Isard, Guy Gur-Ari, Pengcheng Yin, Toju Duke, Anselm
  Levskaya, Sanjay Ghemawat, Sunipa Dev, Henryk Michalewski, Xavier Garcia,
  Vedant Misra, Kevin Robinson, Liam Fedus, Denny Zhou, Daphne Ippolito, David
  Luan, Hyeontaek Lim, Barret Zoph, Alexander Spiridonov, Ryan Sepassi, David
  Dohan, Shivani Agrawal, Mark Omernick, Andrew~M. Dai,
  Thanumalayan~Sankaranarayana Pillai, Marie Pellat, Aitor Lewkowycz, Erica
  Moreira, Rewon Child, Oleksandr Polozov, Katherine Lee, Zongwei Zhou, Xuezhi
  Wang, Brennan Saeta, Mark Diaz, Orhan Firat, Michele Catasta, Jason Wei,
  Kathy Meier-Hellstern, Douglas Eck, Jeff Dean, Slav Petrov, and Noah Fiedel.
  2022.
\newblock \href {https://arxiv.org/abs/2204.02311} {Palm: Scaling language
  modeling with pathways}.

\bibitem[{Clark et~al.(2018)Clark, Cowhey, Etzioni, Khot, Sabharwal, Schoenick,
  and Tafjord}]{clark2018think}
Peter Clark, Isaac Cowhey, Oren Etzioni, Tushar Khot, Ashish Sabharwal, Carissa
  Schoenick, and Oyvind Tafjord. 2018.
\newblock \href {https://arxiv.org/abs/1803.05457} {Think you have solved
  question answering? try arc, the ai2 reasoning challenge}.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and
  Toutanova}]{devlin-etal-2019-bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock {BERT}: Pre-training of deep bidirectional transformers for language
  understanding.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies (NAACL-HLT)}, pages 4171--4186.

\bibitem[{Fujii et~al.(2024)Fujii, Nakamura, Loem, Iida, Ohi, Hattori, Shota,
  Mizuki, Yokota, and Okazaki}]{fujii2024continual}
Kazuki Fujii, Taishi Nakamura, Mengsay Loem, Hiroki Iida, Masanari Ohi, Kakeru
  Hattori, Hirai Shota, Sakae Mizuki, Rio Yokota, and Naoaki Okazaki. 2024.
\newblock Continual pre-training for cross-lingual {LLM} adaptation: Enhancing
  japanese language capabilities.
\newblock In \emph{First Conference on Language Modeling (COLM)}.

\bibitem[{Gao et~al.(2023)Gao, Tow, Abbasi, Biderman, Black, DiPofi, Foster,
  Golding, Hsu, Le~Noac'h, Li, McDonell, Muennighoff, Ociepa, Phang, Reynolds,
  Schoelkopf, Skowron, Sutawika, Tang, Thite, Wang, Wang, and
  Zou}]{eval-harness}
Leo Gao, Jonathan Tow, Baber Abbasi, Stella Biderman, Sid Black, Anthony
  DiPofi, Charles Foster, Laurence Golding, Jeffrey Hsu, Alain Le~Noac'h,
  Haonan Li, Kyle McDonell, Niklas Muennighoff, Chris Ociepa, Jason Phang,
  Laria Reynolds, Hailey Schoelkopf, Aviya Skowron, Lintang Sutawika, Eric
  Tang, Anish Thite, Ben Wang, Kevin Wang, and Andy Zou. 2023.
\newblock A framework for few-shot language model evaluation.

\bibitem[{Grave et~al.(2017)Grave, Joulin, Cissé, Grangier, and
  Jégou}]{ICML-2017-GraveJCGJ}
Edouard Grave, Armand Joulin, Moustapha Cissé, David Grangier, and Hervé
  Jégou. 2017.
\newblock {Efficient softmax approximation for GPUs}.
\newblock In \emph{{Proceedings of the 34th International Conference on Machine
  Learning (ICML)}}, pages 1302--1310.

\bibitem[{Halko et~al.(2011)Halko, Martinsson, and
  Tropp}]{doi:10.1137/090771806}
Nathan Halko, Per-Gunnar Martinsson, and Joel~A. Tropp. 2011.
\newblock \href {https://doi.org/10.1137/090771806} {Finding structure with
  randomness: Probabilistic algorithms for constructing approximate matrix
  decompositions}.
\newblock \emph{SIAM Review}, 53(2):217--288.

\bibitem[{Hoffmann et~al.(2022)Hoffmann, Borgeaud, Mensch, Buchatskaya, Cai,
  Rutherford, de~Las~Casas, Hendricks, Welbl, Clark, Hennigan, Noland,
  Millican, van~den Driessche, Damoc, Guy, Osindero, Simonyan, Elsen, Rae,
  Vinyals, and Sifre}]{Hoffmann2022TrainingCL}
Jordan Hoffmann, Sebastian Borgeaud, Arthur Mensch, Elena Buchatskaya, Trevor
  Cai, Eliza Rutherford, Diego de~Las~Casas, Lisa~Anne Hendricks, Johannes
  Welbl, Aidan Clark, Tom Hennigan, Eric Noland, Katie Millican, George van~den
  Driessche, Bogdan Damoc, Aurelia Guy, Simon Osindero, Karen Simonyan, Erich
  Elsen, Jack~W. Rae, Oriol Vinyals, and L.~Sifre. 2022.
\newblock Training compute-optimal large language models.
\newblock \emph{ArXiv}, abs/2203.15556.

\bibitem[{Jozefowicz et~al.(2016)Jozefowicz, Vinyals, Schuster, Shazeer, and
  Wu}]{45446}
Rafal Jozefowicz, Oriol Vinyals, Mike Schuster, Noam Shazeer, and Yonghui Wu.
  2016.
\newblock \href {https://arxiv.org/pdf/1602.02410.pdf} {Exploring the limits of
  language modeling}.

\bibitem[{Kaplan et~al.(2020)Kaplan, McCandlish, Henighan, Brown, Chess, Child,
  Gray, Radford, Wu, and Amodei}]{kaplan2020scaling}
Jared Kaplan, Sam McCandlish, Tom Henighan, Tom~B. Brown, Benjamin Chess, Rewon
  Child, Scott Gray, Alec Radford, Jeffrey Wu, and Dario Amodei. 2020.
\newblock \href {https://arxiv.org/abs/2001.08361} {Scaling laws for neural
  language models}.

\bibitem[{Kim et~al.(2024)Kim, Choi, and Jeong}]{kim2024efficient}
Seungduk Kim, Seungtaek Choi, and Myeongho Jeong. 2024.
\newblock \href {https://arxiv.org/abs/2402.14714} {Efficient and effective
  vocabulary expansion towards multilingual large language models}.
\newblock \emph{Preprint}, arXiv:2402.14714.

\bibitem[{Kiyono et~al.(2020)Kiyono, Ito, Konno, Morishita, and
  Suzuki}]{kiyono-etal-2020-tohoku}
Shun Kiyono, Takumi Ito, Ryuto Konno, Makoto Morishita, and Jun Suzuki. 2020.
\newblock Tohoku-{AIP}-{NTT} at {WMT} 2020 news translation task.
\newblock In \emph{Proceedings of the Fifth Conference on Machine Translation
  (WMT)}, pages 145--155.

\bibitem[{Kudo(2018)}]{kudo-2018-subword}
Taku Kudo. 2018.
\newblock Subword regularization: Improving neural network translation models
  with multiple subword candidates.
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association
  for Computational Linguistics (ACL)}, pages 66--75.

\bibitem[{Kudo and Richardson(2018)}]{kudo-richardson-2018-sentencepiece}
Taku Kudo and John Richardson. 2018.
\newblock {S}entence{P}iece: A simple and language independent subword
  tokenizer and detokenizer for neural text processing.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 66--71.

\bibitem[{Kurihara et~al.(2022)Kurihara, Kawahara, and
  Shibata}]{kurihara-etal-2022-jglue}
Kentaro Kurihara, Daisuke Kawahara, and Tomohide Shibata. 2022.
\newblock {JGLUE}: {J}apanese general language understanding evaluation.
\newblock In \emph{Proceedings of the Thirteenth Language Resources and
  Evaluation Conference (LREC)}, pages 2957--2966.

\bibitem[{Lan et~al.(2020)Lan, Chen, Goodman, Gimpel, Sharma, and
  Soricut}]{Lan2020ALBERT:}
Zhenzhong Lan, Mingda Chen, Sebastian Goodman, Kevin Gimpel, Piyush Sharma, and
  Radu Soricut. 2020.
\newblock Albert: A lite bert for self-supervised learning of language
  representations.
\newblock In \emph{International Conference on Learning Representations
  (ICLR)}.

\bibitem[{Le~Scao et~al.(2022)Le~Scao, Wang, Hesslow, Bekman, Bari, Biderman,
  Elsahar, Muennighoff, Phang, Press, Raffel, Sanh, Shen, Sutawika, Tae, Yong,
  Launay, and Beltagy}]{le-scao-etal-2022-language}
Teven Le~Scao, Thomas Wang, Daniel Hesslow, Stas Bekman, M~Saiful Bari, Stella
  Biderman, Hady Elsahar, Niklas Muennighoff, Jason Phang, Ofir Press, Colin
  Raffel, Victor Sanh, Sheng Shen, Lintang Sutawika, Jaesung Tae, Zheng~Xin
  Yong, Julien Launay, and Iz~Beltagy. 2022.
\newblock What language model to train if you have one million {GPU} hours?
\newblock In \emph{Findings of the Association for Computational Linguistics:
  EMNLP 2022}, pages 765--782.

\bibitem[{Libovick{\'y} et~al.(2022)Libovick{\'y}, Schmid, and
  Fraser}]{libovicky-etal-2022-dont}
Jind{\v{r}}ich Libovick{\'y}, Helmut Schmid, and Alexander Fraser. 2022.
\newblock Why don{'}t people use character-level machine translation?
\newblock In \emph{Findings of the Association for Computational Linguistics:
  ACL 2022}, pages 2470--2485.

\bibitem[{Lieber et~al.(2021)Lieber, Sharir, Lenz, and Shoham}]{J1WhitePaper}
Opher Lieber, Or~Sharir, Barak Lenz, and Yoav Shoham. 2021.
\newblock Jurassic-1: Technical details and evaluation.
\newblock Technical report, AI21 Labs.

\bibitem[{Merity et~al.(2017)Merity, Xiong, Bradbury, and
  Socher}]{DBLP:journals/corr/MerityXBS16}
Stephen Merity, Caiming Xiong, James Bradbury, and Richard Socher. 2017.
\newblock {Pointer Sentinel Mixture Models}.
\newblock In \emph{Proceedings of the 5th International Conference on Learning
  Representations (ICLR)}.

\bibitem[{Mihaylov et~al.(2018)Mihaylov, Clark, Khot, and
  Sabharwal}]{mihaylov-etal-2018-suit}
Todor Mihaylov, Peter Clark, Tushar Khot, and Ashish Sabharwal. 2018.
\newblock Can a suit of armor conduct electricity? a new dataset for open book
  question answering.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 2381--2391.

\bibitem[{Minixhofer et~al.(2022)Minixhofer, Paischer, and
  Rekabsaz}]{minixhofer-etal-2022-wechsel}
Benjamin Minixhofer, Fabian Paischer, and Navid Rekabsaz. 2022.
\newblock {WECHSEL}: Effective initialization of subword embeddings for
  cross-lingual transfer of monolingual language models.
\newblock In \emph{Proceedings of the 2022 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies (NAACL-HLT)}, pages 3992--4006.

\bibitem[{M{\"u}ller and Laurent(2022)}]{muller2022cedille}
Martin M{\"u}ller and Florian Laurent. 2022.
\newblock \href {https://api.semanticscholar.org/CorpusID:246634768} {Cedille:
  A large autoregressive {F}rench language model}.
\newblock \emph{ArXiv}, abs/2202.03371.

\bibitem[{Radford et~al.(2018{\natexlab{a}})Radford, Narasimhan, Salimans, and
  Sutskever}]{radford2018improving}
Alec Radford, Karthik Narasimhan, Tim Salimans, and Ilya Sutskever.
  2018{\natexlab{a}}.
\newblock Improving language understanding by generative pre-training.

\bibitem[{Radford et~al.(2018{\natexlab{b}})Radford, Wu, Child, Luan, Amodei,
  and Sutskever}]{noauthororeditor}
Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, and Ilya
  Sutskever. 2018{\natexlab{b}}.
\newblock Language models are unsupervised multitask learners.

\bibitem[{Sakaguchi et~al.(2021)Sakaguchi, Bras, Bhagavatula, and
  Choi}]{10.1145/3474381}
Keisuke Sakaguchi, Ronan~Le Bras, Chandra Bhagavatula, and Yejin Choi. 2021.
\newblock Winogrande: an adversarial winograd schema challenge at scale.
\newblock \emph{Communications of the ACM}, 64(9):99–106.

\bibitem[{Sennrich et~al.(2016)Sennrich, Haddow, and
  Birch}]{sennrich-etal-2016-neural}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016.
\newblock Neural machine translation of rare words with subword units.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (ACL)}, pages 1715--1725.

\bibitem[{Shoeybi et~al.(2020)Shoeybi, Patwary, Puri, LeGresley, Casper, and
  Catanzaro}]{shoeybi2020megatronlm}
Mohammad Shoeybi, Mostofa Patwary, Raul Puri, Patrick LeGresley, Jared Casper,
  and Bryan Catanzaro. 2020.
\newblock \href {https://arxiv.org/abs/1909.08053} {Megatron-lm: Training
  multi-billion parameter language models using model parallelism}.

\bibitem[{Soboleva et~al.(2023)Soboleva, Al-Khateeb, Myers, Steeves, Hestness,
  and Dey}]{cerebras2023slimpajama}
Daria Soboleva, Faisal Al-Khateeb, Robert Myers, Jacob~R Steeves, Joel
  Hestness, and Nolan Dey. 2023.
\newblock {SlimPajama: A 627B token cleaned and deduplicated version of
  RedPajama}.

\bibitem[{Takase et~al.(2024)Takase, Kiyono, Kobayashi, and
  Suzuki}]{takase2024spike}
Sho Takase, Shun Kiyono, Sosuke Kobayashi, and Jun Suzuki. 2024.
\newblock \href {https://arxiv.org/abs/2312.16903} {Spike no more: Stabilizing
  the pre-training of large language models}.

\bibitem[{Takase and Kobayashi(2020)}]{NEURIPS2020_275d7fb2}
Sho Takase and Sosuke Kobayashi. 2020.
\newblock All word embeddings from one embedding.
\newblock In \emph{Advances in Neural Information Processing Systems
  (NeurIPS)}, volume~33, pages 3775--3785.

\bibitem[{Takase et~al.(2019)Takase, Suzuki, and Nagata}]{Takase:chara:ngram}
Sho Takase, Jun Suzuki, and Masaaki Nagata. 2019.
\newblock Character n-gram embeddings to improve rnn language models.
\newblock In \emph{Proceedings of the Thirty-Third AAAI Conference on
  Artificial Intelligence (AAAI)}, pages 5074--5082.

\bibitem[{Tikhonov and Ryabinin(2021)}]{tikhonov-ryabinin-2021-heads}
Alexey Tikhonov and Max Ryabinin. 2021.
\newblock {I}t{'}s {A}ll in the {H}eads: {U}sing {A}ttention {H}eads as a
  {B}aseline for {C}ross-{L}ingual {T}ransfer in {C}ommonsense {R}easoning.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  ACL-IJCNLP 2021}, pages 3534--3546.

\bibitem[{Touvron et~al.(2023)Touvron, Martin, Stone, Albert, Almahairi,
  Babaei, Bashlykov, Batra, Bhargava, Bhosale, Bikel, Blecher, Ferrer, Chen,
  Cucurull, Esiobu, Fernandes, Fu, Fu, Fuller, Gao, Goswami, Goyal, Hartshorn,
  Hosseini, Hou, Inan, Kardas, Kerkez, Khabsa, Kloumann, Korenev, Koura,
  Lachaux, Lavril, Lee, Liskovich, Lu, Mao, Martinet, Mihaylov, Mishra,
  Molybog, Nie, Poulton, Reizenstein, Rungta, Saladi, Schelten, Silva, Smith,
  Subramanian, Tan, Tang, Taylor, Williams, Kuan, Xu, Yan, Zarov, Zhang, Fan,
  Kambadur, Narang, Rodriguez, Stojnic, Edunov, and Scialom}]{touvron2023llama}
Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine
  Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale,
  Dan Bikel, Lukas Blecher, Cristian~Canton Ferrer, Moya Chen, Guillem
  Cucurull, David Esiobu, Jude Fernandes, Jeremy Fu, Wenyin Fu, Brian Fuller,
  Cynthia Gao, Vedanuj Goswami, Naman Goyal, Anthony Hartshorn, Saghar
  Hosseini, Rui Hou, Hakan Inan, Marcin Kardas, Viktor Kerkez, Madian Khabsa,
  Isabel Kloumann, Artem Korenev, Punit~Singh Koura, Marie-Anne Lachaux,
  Thibaut Lavril, Jenya Lee, Diana Liskovich, Yinghai Lu, Yuning Mao, Xavier
  Martinet, Todor Mihaylov, Pushkar Mishra, Igor Molybog, Yixin Nie, Andrew
  Poulton, Jeremy Reizenstein, Rashi Rungta, Kalyan Saladi, Alan Schelten, Ruan
  Silva, Eric~Michael Smith, Ranjan Subramanian, Xiaoqing~Ellen Tan, Binh Tang,
  Ross Taylor, Adina Williams, Jian~Xiang Kuan, Puxin Xu, Zheng Yan, Iliyan
  Zarov, Yuchen Zhang, Angela Fan, Melanie Kambadur, Sharan Narang, Aurelien
  Rodriguez, Robert Stojnic, Sergey Edunov, and Thomas Scialom. 2023.
\newblock \href {https://arxiv.org/abs/2307.09288} {Llama 2: Open foundation
  and fine-tuned chat models}.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{NIPS2017_7181}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, \L~ukasz Kaiser, and Illia Polosukhin. 2017.
\newblock Attention is all you need.
\newblock In \emph{Advances in Neural Information Processing Systems 30
  (NIPS)}, pages 5998--6008.

\bibitem[{Wenzek et~al.(2020)Wenzek, Lachaux, Conneau, Chaudhary, Guzm{\'a}n,
  Joulin, and Grave}]{wenzek-etal-2020-ccnet}
Guillaume Wenzek, Marie-Anne Lachaux, Alexis Conneau, Vishrav Chaudhary,
  Francisco Guzm{\'a}n, Armand Joulin, and Edouard Grave. 2020.
\newblock {CCN}et: Extracting high quality monolingual datasets from web crawl
  data.
\newblock In \emph{Proceedings of the Twelfth Language Resources and Evaluation
  Conference (LREC)}, pages 4003--4012.

\bibitem[{Wortsman et~al.(2024)Wortsman, Liu, Xiao, Everett, Alemi, Adlam,
  Co-Reyes, Gur, Kumar, Novak, Pennington, Sohl-Dickstein, Xu, Lee, Gilmer, and
  Kornblith}]{wortsman2024smallscale}
Mitchell Wortsman, Peter~J Liu, Lechao Xiao, Katie~E Everett, Alexander~A
  Alemi, Ben Adlam, John~D Co-Reyes, Izzeddin Gur, Abhishek Kumar, Roman Novak,
  Jeffrey Pennington, Jascha Sohl-Dickstein, Kelvin Xu, Jaehoon Lee, Justin
  Gilmer, and Simon Kornblith. 2024.
\newblock Small-scale proxies for large-scale transformer training
  instabilities.
\newblock In \emph{The Twelfth International Conference on Learning
  Representations (ICLR)}.

\bibitem[{Xiong et~al.(2020)Xiong, Yang, He, Zheng, Zheng, Xing, Zhang, Lan,
  Wang, and Liu}]{DBLP:conf/icml/XiongYHZZXZLWL20}
Ruibin Xiong, Yunchang Yang, Di~He, Kai Zheng, Shuxin Zheng, Chen Xing,
  Huishuai Zhang, Yanyan Lan, Liwei Wang, and Tie{-}Yan Liu. 2020.
\newblock On layer normalization in the transformer architecture.
\newblock In \emph{Proceedings of the 37th International Conference on Machine
  Learning (ICML)}, pages 10524--10533.

\bibitem[{Xue et~al.(2021)Xue, Constant, Roberts, Kale, Al-Rfou, Siddhant,
  Barua, and Raffel}]{xue-etal-2021-mt5}
Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al-Rfou, Aditya
  Siddhant, Aditya Barua, and Colin Raffel. 2021.
\newblock m{T}5: A massively multilingual pre-trained text-to-text transformer.
\newblock In \emph{Proceedings of the 2021 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies (NAACL-HLT)}, pages 483--498.

\bibitem[{Yamada and Ri(2024)}]{yamada2024leia}
Ikuya Yamada and Ryokan Ri. 2024.
\newblock \href {https://arxiv.org/abs/2402.11485} {Leia: Facilitating
  cross-lingual knowledge transfer in language models with entity-based data
  augmentation}.
\newblock \emph{Preprint}, arXiv:2402.11485.

\bibitem[{Yong et~al.(2023)Yong, Schoelkopf, Muennighoff, Aji, Adelani,
  Almubarak, Bari, Sutawika, Kasai, Baruwa, Winata, Biderman, Raff, Radev, and
  Nikoulina}]{yong-etal-2023-bloom}
Zheng~Xin Yong, Hailey Schoelkopf, Niklas Muennighoff, Alham~Fikri Aji,
  David~Ifeoluwa Adelani, Khalid Almubarak, M~Saiful Bari, Lintang Sutawika,
  Jungo Kasai, Ahmed Baruwa, Genta Winata, Stella Biderman, Edward Raff,
  Dragomir Radev, and Vassilina Nikoulina. 2023.
\newblock {BLOOM}+1: Adding language support to {BLOOM} for zero-shot
  prompting.
\newblock In \emph{Proceedings of the 61st Annual Meeting of the Association
  for Computational Linguistics (ACL)}, pages 11682--11703.

\bibitem[{Zellers et~al.(2019)Zellers, Holtzman, Bisk, Farhadi, and
  Choi}]{zellers-etal-2019-hellaswag}
Rowan Zellers, Ari Holtzman, Yonatan Bisk, Ali Farhadi, and Yejin Choi. 2019.
\newblock {H}ella{S}wag: Can a machine really finish your sentence?
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics (ACL)}, pages 4791--4800.

\bibitem[{Zhang et~al.(2022)Zhang, Roller, Goyal, Artetxe, Chen, Chen, Dewan,
  Diab, Li, Lin, Mihaylov, Ott, Shleifer, Shuster, Simig, Koura, Sridhar, Wang,
  and Zettlemoyer}]{zhang2022opt}
Susan Zhang, Stephen Roller, Naman Goyal, Mikel Artetxe, Moya Chen, Shuohui
  Chen, Christopher Dewan, Mona Diab, Xian Li, Xi~Victoria Lin, Todor Mihaylov,
  Myle Ott, Sam Shleifer, Kurt Shuster, Daniel Simig, Punit~Singh Koura, Anjali
  Sridhar, Tianlu Wang, and Luke Zettlemoyer. 2022.
\newblock \href {https://arxiv.org/abs/2205.01068} {Opt: Open pre-trained
  transformer language models}.

\end{thebibliography}

\documentclass[11pt]{article}

\usepackage{acl}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{inconsolata}

\usepackage{graphicx}

\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{mathtools}
\usepackage{amsthm}

\title{Large Vocabulary Size Improves Large Language Models}

\author{Sho Takase \hspace{1.5em} Ryokan Ri \hspace{1.5em} Shun Kiyono \hspace{1.5em} Takuya Kato \\
  SB Intuitions \\
  {\tt \{sho.takase, ryokan.ri, shun.kiyono, takuya.kato\}@sbintuitions.co.jp} \\
  }

\begin{document}
\maketitle
\begin{abstract}
This paper empirically investigates the relationship between subword vocabulary size and the performance of large language models (LLMs) to provide insights on how to define the vocabulary size.
Experimental results show that larger vocabulary sizes lead to better performance in LLMs.
Moreover, we consider a continual training scenario where a pre-trained language model is trained on a different target language.
We introduce a simple method to use a new vocabulary instead of the pre-defined one.
We show that using the new vocabulary outperforms the model with the vocabulary used in pre-training.
\end{abstract}

\section{Introduction}


Since the GPT series demonstrated that Large Language Models (LLMs) excel in complex reasoning tasks~\cite{radford2018improving,noauthororeditor,NEURIPS2020_1457c0d6}, they have rapidly become indispensable tools for various natural language processing tasks.
To construct better LLMs, previous studies have addressed theoretical analyses of internal layers~\cite{DBLP:conf/icml/XiongYHZZXZLWL20,takase2024spike} and conducted extensive experiments to provide empirical findings~\cite{kaplan2020scaling,Hoffmann2022TrainingCL,wortsman2024smallscale}.
For example, \citet{Hoffmann2022TrainingCL} reported the compute-optimal training configuration, which determines suitable parameter and training data sizes for a given computational resource.

In contrast, although previous studies have explored the properties of internal layers in LLMs, parameters related to the vocabulary, the embedding and output layers, are under-explored.
Specifically, there are no well-established findings on how to determine the subword vocabulary size, which defines the parameter size of the embedding and output layers.
As a standard strategy, a vocabulary size in the 30k-60k range is used for monolingual LLMs~\cite{noauthororeditor,NEURIPS2020_1457c0d6,black2022gptneox20b,zhang2022opt,touvron2023llama}, while around 250k is used for multilingual LLMs~\cite{chowdhery2022palm,le-scao-etal-2022-language}.
For monolingual LLMs, a larger vocabulary size has been discussed in terms of efficiency during the inference phase~\cite{almazrouei2023falcon}.
However, the question remains: does a larger vocabulary size offer any advantages for the quality of monolingual LLMs?
To address this question, we empirically investigate the relationship between vocabulary size and performance on downstream tasks.

We conduct experiments on two languages: English, which is widely used, and Japanese, which is character-rich.
We show that a larger vocabulary size improves the performance of LLMs in both languages.
In addition to training from scratch, we consider the continual training scenario.
When adapting a pre-trained LLM to another language, it may be beneficial to reconstruct an appropriate vocabulary instead of reusing the original vocabulary.
For this purpose, we propose a strategy to swap parameters related to the vocabulary.
We demonstrate that using the reconstructed vocabulary can improve performance.

\section{Vocabulary Construction}
\label{sec:vocab_construction}
To construct subword vocabularies, there are two widely used algorithms: Byte-Pair Encoding (BPE)~\cite{sennrich-etal-2016-neural} and unigram language model~\cite{kudo-2018-subword}.
In this study, we use the unigram language model implemented in SentencePiece~\cite{kudo-richardson-2018-sentencepiece}.
For each language, we use the following vocabulary sizes: 5k, 10k, 50k, 100k and 500k.


We conduct experiments on two languages: English and Japanese.
For the English training data, we extract English corpora from SlimPajama~\cite{cerebras2023slimpajama}, excluding the book corpus, which was reported to have copyright infringement issues.
For the Japanese training data, we extract the Japanese portion of CommonCrawl corpus with the language identification and document deduplication applied using CCNet~\cite{wenzek-etal-2020-ccnet}.
For the vocabulary construction, we sample a small portion (50GB) from each language training data.

\section{Experiments on Vocabulary Size}
\label{sec:main_exp}

\subsection{Settings}
\label{sec:main_setting}

To investigate the relationship between vocabulary size and performance, we train Transformer-based language models on the training data described in Section \ref{sec:vocab_construction}.
Table \ref{tab:datasize} shows the number of tokens in the training data calculated from each vocabulary set.
As shown in this table, the number of tokens varies drastically based on the vocabulary size.
Therefore, we must take care not to give any unfair advantages to any setting.

For example, with a fixed number of training tokens, the 500k vocabulary model trains for around 1.5 epochs in English and 2 epochs in Japanese, while the 5k vocabulary model trains for only 1 epoch.
The larger vocabulary size has an advantage of seeing more data in this configuration.
In contrast, with a fixed number of training epochs, the 5k vocabulary model consumes much more computational resources than the larger vocabulary models.
Especially in Japanese, where the 5k vocabulary model contains about twice as much tokens as the 500k vocabulary model in 1 epoch.
Because the performance of LLMs is correlated with the computational costs during training~\cite{kaplan2020scaling}, this configuration might favor smaller vocabulary sizes.
Thus, we prepare two training configurations: 1T tokens and 1 epoch\footnote{In addition to the training data size, we have to discuss the number of parameters for a fair comparison because the model with the small vocabulary size contains less parameters for the embedding and output layers. However, as described in Appendix \ref{sec:comp_paramsize}, the model with the small vocabulary size does not improve the performance when we increase the number of parameters related to the vocabulary. Thus, we focus only on varying the training data size in our main experiments.}.

For hyper-parameters of the language model, we use the GPT-3 Large setting described in \citet{NEURIPS2020_1457c0d6}.
We set the number of layers $24$ and the hidden dimension size $1536$.
In this setting, the number of parameters for internal layers is 680M.
We use Megatron-LM~\cite{shoeybi2020megatronlm}\footnote{\href{https://github.com/NVIDIA/Megatron-LM}{https://github.com/NVIDIA/Megatron-LM}} as our codebase to train large language models.
To stabilize the training, we use the scaled embed technique~\cite{takase2024spike}.

We evaluate each model on the commonsense reasoning tasks.
For English, we use PIQA~\cite{Bisk_Zellers_Lebras_Gao_Choi_2020}, OpenBookQA (OBQA)~\cite{mihaylov-etal-2018-suit}, HellaSwag~\cite{zellers-etal-2019-hellaswag}, WinoGrande~\cite{10.1145/3474381} and ARC easy and challenge~\cite{clark2018think}.
For Japanese, we use JSQuAD and JCommonsenseQA (JCQA) from JGLUE~\cite{kurihara-etal-2022-jglue}, the Japanese portion of XWinograd~\cite{tikhonov-ryabinin-2021-heads}, and JAQKET\footnote{\href{https://sites.google.com/view/project-aio/competition1}{https://sites.google.com/view/project-aio/competition1}}.
Following the previous study~\cite{touvron2023llama}, we use the normalized likelihood in evaluation~\cite{NEURIPS2020_1457c0d6,eval-harness}.

\begin{table}[!t]
  \centering
  \footnotesize
  \begin{tabular}{ l | c c } \hline
  \multicolumn{1}{c|}{\#Vocab} & English & Japanese \\ \hline \hline
  5k & 830B & 950B \\
  10k & 750B & 750B \\
  50k & 670B & 590B \\
  100k & 650B & 550B \\
  500k & 640B & 490B \\ \hline
  \end{tabular}
  \caption{The number of tokens in training data tokenized by each vocabulary.}
  \label{tab:datasize}
\end{table}

\subsection{Results}

\begin{table*}[!t]
  \centering
  \footnotesize
  \begin{tabular}{ l | c c c c c c | c } \hline
  \multicolumn{1}{c|}{\#Vocab} & PIQA & OBQA & HellaSwag & WinoGrande & ARC-e & ARC-c & Avg.\\ \hline \hline
  \multicolumn{8}{c}{1T tokens} \\ \hline \hline
  5k & 69.9 & 33.2 & 51.0 & 55.2 & 49.6 & 27.7 & 47.8 ($\pm 0.0$)\\
  10k & 71.2 & 33.4 & 51.5 & 55.2 & 50.6 & 27.1 & 48.2 ($+ 0.4$) \\
  50k & \textbf{71.7} & 32.8 & 53.9 & 54.5 & 50.8 & 27.7 & 48.6 ($+0.8$)\\
  100k & 70.9 & 33.4 & 53.9 & 54.8 & 54.3 & 27.7 & 49.2 ($+1.4$)\\
  500k & 71.4 & \textbf{34.0} & \textbf{55.3} & \textbf{57.5} & \textbf{55.1} & \textbf{28.3} & \textbf{50.3} ($+2.5$)\\ \hline
  \multicolumn{8}{c}{1 Epoch} \\ \hline \hline  
  5k & 70.1 & 32.4 & 50.9 & \textbf{55.2} & 50.2 & \textbf{28.5} & 47.9 ($\pm 0.0$) \\
  10k & 71.1 & 33.6 & 50.6 & 55.7 & 49.0 & 27.1 & 47.9 ($\pm 0.0$) \\
  50k & 70.6 & 33.6 & 52.1 & 53.8 & 52.3 & 27.3 & 48.3 ($+ 0.4$)\\
  100k & \textbf{71.7} & 33.8 & 53.4 & 54.7 & 52.7 & 27.6 & 49.0 ($+ 1.1$)\\
  500k & 70.4 & \textbf{34.2} & \textbf{54.3} & 55.1 & \textbf{54.0} & 28.2 & \textbf{49.4} ($+ 1.5$)\\ \hline
  \end{tabular}
  \caption{The performance on English commonsense reasoning tasks in training 1T tokens and 1 epoch.}
  \label{tab:main_result_en}
\end{table*}

\begin{table*}[!t]
  \centering
  \footnotesize
  \begin{tabular}{ l | c c c c | c } \hline
  \multicolumn{1}{c|}{\#Vocab} & JSQuAD & JCQA & XWinograd & JAQKET & Avg. \\ \hline \hline
  \multicolumn{6}{c}{1T tokens} \\ \hline \hline
  5k & 58.1 & 68.1 & 58.9 & 12.5 & 49.4 ($\pm 0.0$) \\
  10k & 61.2 & 67.2 & 59.0 & 23.3 & 52.7 ($+ 3.3$) \\
  50k & 61.8 & 71.6 &  59.0 & 29.2 & 55.4 ($+ 6.0$) \\
  100k & 62.1 & \textbf{71.9} & \textbf{59.6} & 34.9 & 57.1 ($+ 7.7$) \\
  500k & \textbf{64.5} & 71.6 & 59.3 & \textbf{38.9} & \textbf{58.6} ($+ 9.2$) \\ \hline
  \multicolumn{6}{c}{1 Epoch} \\ \hline \hline
  5k & 57.7 & 68.1 & 58.8 & 14.4 & 49.8 ($\pm 0.0$) \\
  10k & 57.7 & 63.4 & \textbf{60.0} & 22.0 & 50.8 ($+ 1.0$) \\
  50k & 60.9 & 69.1 & 58.5 & 28.7 & 54.3 ($+ 4.5$) \\
  100k & 61.3 & \textbf{70.1} & 58.7 & 31.0 & 55.3 ($+ 5.5$) \\
  500k & \textbf{63.2} & 69.8 & 57.7 & \textbf{34.1} & \textbf{56.2} ($+ 6.4$) \\ \hline
  \end{tabular}
  \caption{The performance on Japanese commonsense reasoning tasks in training 1T tokens and 1 epoch.}
  \label{tab:main_result_ja}
\end{table*}



Tables \ref{tab:main_result_en} and \ref{tab:main_result_ja} present the performance of the models trained with 1T tokens and 1 epoch.
For each configuration, we show the average score of each task, and the improvement of the average score from the 5k vocabulary model.

As shown by the average scores, for both English and Japanese, larger vocabulary sizes lead to better performance.
The improvement is particularly notable in Japanese, largely due to the gains in JAQKET.
Unlike the other tasks where the model selects answers from provided candidates, JAQKET is a factoid QA task where the model generates answers without any candidates.
This suggests that a larger vocabulary size particularly benefits generation tasks.

In addition, the larger vocabulary size achieves better performance in either situation where we fix the number of training tokens or training epochs.
With a fixed number of epochs, the larger vocabulary size settings, e.g., 100k and 500k, use a much smaller number of training tokens (Table \ref{tab:datasize}).
This means that the larger vocabulary size also improves the training efficiency because we can obtain a better model with a smaller computational cost.
In fact, the GPU hours\footnote{We used A100 80GB for all experiments.} in the 100k setting are 0.7 times shorter than in the 5k when we fix the number of training epochs in Japanese\footnote{Since the larger vocabulary size slows the computation of the output distribution, we should use an efficient way such as the adaptive softmax~\cite{ICML-2017-GraveJCGJ} in practice.}.

\section{Experiments on Continual Training}
\label{sec:exp_continual_training}

\begin{table*}[!t]
  \vspace{-3.5mm}
  \centering
  \footnotesize
  \begin{tabular}{ l l | c c c c | l } \hline
  Setting & \multicolumn{1}{c|}{\#Vocab} & JSQuAD & JCQA & XWinograd & JAQKET & \multicolumn{1}{c}{Avg.} \\ \hline \hline
  From scratch & 100k & 71.8 & 76.0 & 63.6 & 54.2 & 66.4 \\ \hline
  Llama2 (w/o train) & 32k & 71.2 & 60.8 & 62.4 & 15.3 & 52.4 ($\pm 0.0$) \\
  Llama2 vocab & 32k & 80.7 & 79.4 & \textbf{72.6} & 47.7 & 70.1 ($+ 17.7$)\\
  Swap & 100k & 79.2 & \textbf{80.2} & 67.5 & 56.3 & 70.8 ($+18.4$)\\
  Swap\&Insert & 100k & \textbf{81.9} & \textbf{80.2} & 69.2 & \textbf{61.2} & \textbf{73.1} ($+ 20.7$)\\ \hline
  \citet{fujii2024continual} & 100k & 81.6 & 77.6 & 69.1 & 61.1 & 72.4 ($+ 20.0$) \\ \hline
  \end{tabular}
  \caption{The performance on Japanese commonsense reasoning tasks in the continual training from Llama2.}
  \label{tab:continual_training}
\end{table*}

\subsection{Increasing Vocabulary Size}
\label{sec:swap_method}
Section \ref{sec:main_exp} shows that the larger vocabulary size is useful in constructing LLMs from scratch.
In contrast, nowadays, we often start from a high-quality pre-trained model such as the Llama series~\cite{touvron2023llama} and continue training on the target language data~\cite{muller2022cedille,yong-etal-2023-bloom,yamada2024leia}.

Here, we check if we can readily increase the vocabulary size from the pre-trained model.
Similar techniques have been explored as vocabulary expansion~\cite{fujii2024continual,kim2024efficient} or sophisticated embedding initialization using cross-lingual word embeddings~\cite{minixhofer-etal-2022-wechsel}, but our focus here is to check if we could increase the vocabulary size in a rather simplistic way.
We consider a situation where we construct an entirely new vocabulary independently of the original vocabulary.

Let $V_{orig}$ and $V_{new}$ be the vocabulary set of the pre-trained model and a newly constructed vocabulary set respectively, and let $d$ be the dimension size of each layer.
To exploit knowledge learned in the pre-trained embedding matrix, we construct a new embedding matrix $E_{new} \in \mathbb{R}^{|V_{new}|\times d}$ from the original embedding matrix $E_{orig} \in \mathbb{R}^{|V_{orig}|\times d}$ with the way inspired by the randomized algorithm~\cite{doi:10.1137/090771806}:

\begin{align}
E_{new} = \frac{W E_{orig}}{\sqrt{|V_{orig}|}}, \label{eq:linear}
\end{align}
where $W \in \mathbb{R}^{|V_{new}| \times |V_{orig}|}$ is the random matrix whose elements are sampled from the standard normal distribution independently.
To maintain the standard deviation of $E_{orig}$ in $E_{new}$, we scale the matrix multiplication by $\frac{1}{\sqrt{|V_{orig}|}}$\footnote{We assume that $E_{orig}$ contains independent random variables with mean $0$ and variance $\mathrm{var}(E_{orig})$. Then, the variance of the matrix multiplication $W E_{orig}$ has mean $0$ and variance $\mathrm{var}(E_{orig}) \times |V_{orig}|$.}.

In the naive way, we swap $E_{new}$ with $E_{orig}$.
However, Equation \ref{eq:linear} randomizes embeddings even if $V_{new}$ contains the corresponding subwords which may possess useful knowledge transferable to the new model.
Therefore, we insert the pre-trained embedding in $E_{orig}$ into $E_{new}$ if the corresponding subword is included in both $V_{orig}$ and $V_{new}$\footnote{See Appendix \ref{sec:explanation_on_insert} for more details.}.
For the output layer, we construct a new weight matrix with the same manner.

\subsection{Results}
\label{sec:results_on_continual_training}
We train the Llama2 7B parameter model~\cite{touvron2023llama} with 100B tokens on our Japanese training data.
We use the Japanese vocabulary whose size is 100k.
Table \ref{tab:continual_training} shows results on Japanese commonsense reasoning tasks.
In this table, `Swap' uses new parameters related to the vocabulary without inserting the corresponding pre-trained parameters.
We train a language model from scratch to compare the effectiveness of the continual training.
Moreover, we compare the embedding initialization method by \citet{fujii2024continual} because their study is the same situation: continual training of Llama2 on Japanese data.

Table \ref{tab:continual_training} shows that `Swap' and `Swap\&Insert' outperform the model using the original Llama2 vocabulary even though these settings randomize parameters related to the vocabulary.
This result indicates that it is better to prepare an appropriate vocabulary even in the continual training situation.
Moreover, the insertion strategy achieves further improvement.
The `Swap\&Insert' outperforms the method of \citet{fujii2024continual}, which initializes an embedding of the new subword with the average of the pre-trained embeddings\footnote{For the existing subwords, their method uses the pre-trained embeddings. Thus, their method is regarded as using the `Insert' strategy.}, and thus, the `Swap\&Insert' is simple but effective.


\section{Related Work}
Before the paradigm of subword units and LLMs, researchers sometimes needed to handle the large vocabulary size such as more than 100k to decrease the number of unknown words.
For example, the vocabulary sizes of One Billion Word Benchmark and WikiText-103 are about 800k and 300k respectively~\cite{41880,DBLP:journals/corr/MerityXBS16}.
Some previous studies reported that character-level information was useful for neural language models with the large vocabulary size~\cite{45446,Takase:chara:ngram}.
In this paradigm, \citet{chen-etal-2019-large} explored the impact of the vocabulary size.

Since the use of subword units is proposed~\cite{sennrich-etal-2016-neural,kudo-2018-subword}, the vocabulary sizes 30k-60k are widely used as the magic numbers~\citep{libovicky-etal-2022-dont}.
As examples, the BERT and GPT papers use 30k and 40k for their vocabulary sizes respectively without any justification~\cite{NIPS2017_7181,devlin-etal-2019-bert,radford2018improving}.
\citet{kiyono-etal-2020-tohoku} investigated the relation between the performance and the vocabulary size but the maximum vocabulary size of their investigation is too small, i.e., 32k.

For large language models, the vocabulary sizes 30k-60k are also frequently used~\cite{noauthororeditor,touvron2023llama}.
In using the large vocabulary size, the authors claim to support multilinguality~\cite{le-scao-etal-2022-language,xue-etal-2021-mt5} or improve the efficiency~\cite{J1WhitePaper,llama3modelcard}.
In contrast, we investigate the relation between the vocabulary size and the performance of monolingual LLMs on each task.

\section{Conclusion}
In this paper, we empirically investigate the performance of monolingual LLMs when we vary the vocabulary size.
We conduct experiments on two languages: English and Japanese.
Experimental results show that the larger vocabulary size is, the better performance the language model achieves in both languages.
Moreover, we introduce a method to use the entirely new vocabulary in the continual training situation.
We show that using the appropriate vocabulary also improves the performance in the continual training.

\section*{Limitations}
In this study, we conducted experiments on two languages: English and Japanese.
We believe that our findings can be applied to other languages because we do not depend on linguistic features in the subword vocabulary construction.
However, we also agree that it is better to conduct exhaustive experiments on various languages to confirm the generality of our findings.

In this study, we used 500k as the maximum vocabulary size.
Because it is impractical to construct a much larger vocabulary than 500k, we could not investigate the improvement by the tremendously large vocabulary size such as one million and the upper bound of the performance.
The computational time of the vocabulary construction depends on the corpus size and the desired vocabulary size.
We roughly estimate that the vocabulary whose size is larger than one million requires at least over a month in its construction in our environment.

Furthermore, the parameter sizes of internal layers are 680M in training from scratch, and 7B in the continual training.
We consider that the discussions on subword vocabulary size are orthogonal to the parameter size of internal layers, but we would conduct additional experiments with more than 10B parameters if we had a large amount of computational resources.

\bibliographystyle{acl_natbib}


\clearpage

\onecolumn

\appendix

\section{Hyper-parameters}
\label{sec:hyperparam}

Table \ref{tab:hyper_params} shows hyper-parameters used in our main experiments described in Section \ref{sec:main_exp}.

\begin{table}[!t]
\centering
\footnotesize
\begin{tabular}{l | r} \hline
Hyper-parameter & Value \\ \hline \hline
Number of layers & 24 \\
Hidden dimension size & 1536 \\
Number of attention heads & 16 \\
Sequence length & 2048 \\
Batch size & 2048 \\
Learning rate & 3e-4 \\
Learning rate scheduler & Cosine \\
Warmup ratio & 0.01 \\
Adam $\beta_1$ & 0.9 \\
Adam $\beta_2$ & 0.95 \\
Weight decay & 0.01 \\
Gradient clipping & 1.0 \\ \hline
\end{tabular}
  \caption{Hyper-parameters used in experiments described in Section \ref{sec:main_exp}.}
  \label{tab:hyper_params}
\end{table}


\section{Formula of `Insert' in Section \ref{sec:swap_method}}
\label{sec:explanation_on_insert}
We formulate the procedure of `Insert' in Section \ref{sec:swap_method}.
Let $e_{i}^{orig}$ and $e_{i}^{new}$ be the $i$-th row vectors of $E_{orig}$ and $E_{new}$, and let $w_{i}^{orig}$ and $w_{i}^{new}$ be the corresponding subwords to $e_{i}^{orig}$ and $e_{i}^{new}$.
The `Insert' function, $\mathrm{Insert(\cdot)}$,  replaces $e_{i}^{new}$ with $e_{i}^{orig}$ when the corresponding subword is included in the original vocabulary $V_{orig}$ as follows:
\begin{align}
\mathrm{Insert}(e_{i}^{new}) = 
  \begin{cases}
  e_{j}^{orig} & \text{if $w_{i}^{new} \in V_{orig} \land w_{j}^{orig} = w_{i}^{new}$} \\
  e_{i}^{new} & \text{otherwise}
  \end{cases}
\end{align}
Therefore, the matrix contains both the randomized embeddings and the original pre-trained embeddings after the `Insert' procedure.
As shown in Section \ref{sec:results_on_continual_training}, this procedure leads to further improvement.

\section{Comparison on Vocabulary Expansion in Continual Training}
\label{sec:comp_vocab_expansion}

\begin{table*}[!t]
  \centering
  \footnotesize
  \begin{tabular}{ l l | c c c c | l } \hline
  Initialization & Vocab type & JSQuAD & JCQA & XWinograd & JAQKET & \multicolumn{1}{c}{Avg.} \\ \hline \hline
  \citet{fujii2024continual} & Expansion & 78.8 & 63.5 & 63.6 & 50.1 & 64.0 \\
  Swap\&Insert & Expansion & 80.7 & 60.6 & 67.4 & 55.9 & 66.2 \\
  \citet{fujii2024continual} & Appropriate & 81.6 & 77.6 & 69.1 & 61.1 & 72.4 \\
  Swap\&Insert & Appropriate & \textbf{81.9} & \textbf{80.2} & \textbf{69.2} & \textbf{61.2} & \textbf{73.1}\\ \hline
  \end{tabular}
  \caption{The performance on Japanese commonsense reasoning tasks in the continual training from Llama2 when we construct the 100k vocabulary with the vocabulary expansion approach and construct the appropriate 100k vocabulary to the Japanese training data.}
  \label{tab:comparison_vocab_expansion_appropriate}
\end{table*}

\begin{table}[!t]
  \centering
  \footnotesize
  \begin{tabular}{l | r} \hline
  \multicolumn{1}{c|}{Type} & Number \\ \hline \hline
  UTF-8 byte pieces & 256 \\
  Alphabet \& number (e.g., a, the, 1) & 5349 \\
  Symbol (e.g., +, =, \#\#) & 209 \\
  Others such as Japanese characters & 1083 \\ \hline \hline
  Total & 6897 \\ \hline
  \end{tabular}
  \caption{The type and number of shared subword units between the original Llama2 vocabulary and appropriate vocabulary, whose size is 100k, to the Japanese data in the continual training.}
  \label{tab:share_vocab_type}
\end{table}

In Section \ref{sec:exp_continual_training}, we conduct the continual training experiment on the scenario where we construct an appropriate vocabulary to the target language.
In this scenario, most subword units in the original vocabulary might be removed.
In contrast, the vocabulary expansion approach maintains the whole original vocabulary because it only adds new subword units to the original vocabulary~\cite{fujii2024continual}.
We investigate which approach is empirically better in this section.

We construct 100k vocabulary with the vocabulary expansion approach, and compare it with the appropriate vocabulary used in Section \ref{sec:exp_continual_training}.
We apply two strategies to initialize the embedding matrix: \citet{fujii2024continual} and our `Swap\&Insert'.
Table \ref{tab:comparison_vocab_expansion_appropriate} shows results of the continual training from Llama2.
This table indicates that using appropriate vocabulary outperforms the vocabulary expansion approach.
The appropriate vocabulary contains more subword units of the target language.
We consider that this property improves the performance.

Table \ref{tab:share_vocab_type} shows the shared subword units between the original Llama2 vocabulary and the appropriate vocabulary.
This table indicates that the number of shared subword units is only about 7000, which is about one-fifth of the original vocabulary.
Moreover, this table suggests that the original vocabulary contains few Japanese subword units because the number of the shared Japanese characters is about 1000.
Therefore, it is better to construct an entirely new vocabulary that is appropriate to the target language.

For the embedding initialization methods, Table \ref{tab:comparison_vocab_expansion_appropriate} shows that our `Swap\&Insert' achieves better averaged score than the method of \citet{fujii2024continual} in the same as the results in Section \ref{sec:exp_continual_training}.
Thus, our approach is also more suitable in the vocabulary expansion situation.

\section{Comparison on Parameter Size}
\label{sec:comp_paramsize}
The smaller vocabulary size lessens the parameter sizes related to the vocabulary in comparison with the larger vocabulary size.
Thus, the smaller vocabulary size might have the disadvantage in the number of parameters.
To confirm this point, we increase the parameters related to the vocabulary for the 5k setting.
Concretely, we expand the dimension of the embedding and output layers, and then modify the dimension size by the linear transformation such as the matrix factorization technique~\cite{Lan2020ALBERT:}\footnote{In contrast, we can reduce the number of parameters for the larger vocabulary size with the matrix factorization technique or more sophisticated way~\cite{NEURIPS2020_275d7fb2}, but we regard the 5k as the baseline in this experiment.}.
Let $|V|$ be the vocabulary size, $d_e$ be the dimension size of the embedding and output layers, and $d$ be the hidden dimension size.
We prepare the expanded embedding layer $E \in \mathbb{R}^{|V| \times d_e}$ and the trainable weight matrix $W \in \mathbb{R}^{d_e \times d}$.
We convert the dimension size of $E$ with the matrix multiplication $E W$.
For the output layer, we convert the dimension with the same manner.
We adjust $d_e = 30720$ for a fair comparison with the 100k setting in terms of the number of parameters.
For other hyper-parameters, we use the values shown in Table \ref{tab:hyper_params}.

Table \ref{tab:result_on_parameter_size} shows the performance on Japanese commonsense reasoning tasks.
This table indicates that the 5k with the expansion does not improve the average score although it increases the number of parameters.
This result suggests that the increase of the parameter size related to the vocabulary has no positive influence on the performance.
In contrast, the 100k achieves much better average score in the similar parameter size.
Therefore, the improvement by the increase of the vocabulary size is orthogonal to the increase of the parameter size.

\begin{table*}[!t]
  \centering
  \footnotesize
  \begin{tabular}{ l | c c | c c c c | c } \hline
  \multicolumn{1}{c|}{\#Vocab} & \multicolumn{1}{c}{Vocab \#Params.} & \multicolumn{1}{c|}{Total \#Params.} & JSQuAD & JCQA & XWinograd & JAQKET & Avg. \\ \hline \hline
  5k & \ \ \ \ 8M & 690M & 58.1 & 68.1 & 58.9 & 12.5 & 49.4 ($\pm 0.0$) \\
  5k w/ Expansion & 200M & 880M & 61.0 & 60.3 & 59.5 & 15.8 & 49.2 ($- 0.2$)\\
  100k & 150M & 840M & \textbf{62.1} & \textbf{71.9} & \textbf{59.6} & \textbf{34.9} & \textbf{57.1} ($+ 7.7$) \\ \hline
  \end{tabular}
  \caption{The performance on Japanese commonsense reasoning tasks when we use 1T tokens for training. For a fair comparison between 5k and 100k, we increase the parameter sizes of the embedding and output layers (Vocab \#Params. in this Table) for 5k with the matrix factorization technique~\cite{Lan2020ALBERT:}.}
  \label{tab:result_on_parameter_size}
\end{table*}

\section{Experiments on Each Training Data Size}
In addition to the 1T tokens in Section \ref{sec:main_exp}, we investigate the performance in other training data sizes: 10B, 50B, 100B, 200B, and 500B tokens.
Tables \ref{tab:each_data_size_en} and \ref{tab:each_data_size_ja} show the results of English and Japanese models when we use each training data size.
These tables show that larger vocabulary sizes lead to better performance for both English and Japanese in all training data sizes in the same as the results in Section \ref{sec:main_exp}.
These tables indicate that our findings are independent from the amount of training data.

The difference of the performance among vocabulary sizes is smaller in the 10B tokens than ones in other training data sizes.
Thus, the small training data size decreases the advantage of the large vocabulary sizes.
These results explain the relation between our findings and the previous study~\cite{ali-etal-2024-tokenizer}.
\citet{ali-etal-2024-tokenizer} concluded that the small vocabulary size such as 30k is sufficient for English monolingual LLMs.
We consider that they led the contrary conclusion to our findings because their training data, which is about 50B tokens, is much smaller than ours.

\begin{table*}[!t]
  \centering
  \footnotesize
  \begin{tabular}{ l | c c c c c c | c } \hline
  \multicolumn{1}{c|}{\#Vocab} & PIQA & OBQA & HellaSwag & WinoGrande & ARC-e & ARC-c & Avg.\\ \hline \hline
  \multicolumn{8}{c}{10B tokens} \\ \hline \hline
  5k & 58.4 & 25.4 & 29.3 & 51.9 & 34.3 & 22.3 & 36.9 ($\pm 0.0$) \\
  10k & 59.1 & 27.8 & 29.6 & \textbf{53.2} & 35.0 & 21.6 & 37.7 ($+ 0.8$) \\
  50k & 62.1 & 26.2 & 29.5 & 49.9 & 38.7 & 21.9 & 38.0 ($+ 1.1$) \\
  100k & \textbf{62.2} & \textbf{27.8} & 29.7 & 49.6 & \textbf{39.0} & 22.7 & 38.5 ($+ 1.6$) \\
  500k & 62.1 & 27.6 & \textbf{30.1} & 51.3 & 38.7 & \textbf{22.9} & \textbf{38.8} ($+ 1.9$) \\ \hline \hline
  \multicolumn{8}{c}{50B tokens} \\ \hline \hline
  5k & 66.7 & 28.0 & 39.0 & \textbf{52.3} & 41.9 & 23.8 & 41.9 ($\pm 0.0$) \\
  10k & 66.3 & 30.4 & 39.5 & 51.0 & 42.6 & 25.3 & 42.5 ($+ 0.6$) \\
  50k & 68.1 & 29.4 & 40.9 & 50.9 & 46.9 & 25.5 & 43.6 ($+ 1.7$) \\
  100k & 68.1 & 31.6 & 42.0 & 51.3 & 46.9 & 25.5 & 44.2 ($+ 2.3$) \\
  500k & \textbf{68.8} & \textbf{32.2} & \textbf{43.1} & 52.0 & \textbf{47.9} & \textbf{25.7} & \textbf{44.9} ($+ 3.0$) \\ \hline \hline
  \multicolumn{8}{c}{100B tokens} \\ \hline \hline
  5k & 67.2 & 30.8 & 42.7 & 52.2 & 44.1 & 26.7 & 44.0 ($\pm 0.0$) \\
  10k & 68.9 & \textbf{31.6} & 42.7 & 51.6 & 45.1 & 25.7 & 44.3 ($+ 0.3$) \\
  50k & 68.9 & 30.8 & 45.1 & 52.6 & 49.1 & 26.2 & 45.5 ($+ 1.5$)\\
  100k & 70.2 & \textbf{31.6} & 46.1 & 52.9 & 49.1 & 25.8 & 45.9 ($+ 1.9$) \\
  500k & \textbf{70.4} & \textbf{31.6} & \textbf{47.0} & \textbf{53.0} & \textbf{50.0} & \textbf{28.2} & \textbf{46.7} ($+ 2.7$) \\ \hline \hline
  \multicolumn{8}{c}{200B tokens} \\ \hline \hline
  5k & 68.8 & 32.8 & 45.3 & 53.4 & 46.0 & 25.3 & 45.2 ($\pm 0.0$) \\
  10k & 69.0 & 31.6 & 46.2 & 53.3 & 45.7 & 26.5 & 45.4 ($+ 0.2$)\\
  50k & 70.5 & 31.0 & 47.9 & 53.8 & 50.0 & 26.1 & 46.6 ($+ 1.4$)\\
  100k & 70.5 & \textbf{33.6} & 49.2 & \textbf{54.6} & 50.6 & 26.2 & 47.4 ($+ 2.2$)\\
  500k & \textbf{70.7} & 33.4 & \textbf{50.2} & 54.3 & \textbf{51.8} & \textbf{29.6} & \textbf{48.3} ($+ 3.1$) \\ \hline \hline
  \multicolumn{8}{c}{500B tokens} \\ \hline \hline
  5k & 69.7 & 32.6 & 49.6 & 52.9 & 47.7 & 26.4 & 46.5 ($\pm 0.0$) \\
  10k & 70.8 & \textbf{34.2} & 49.7 & 54.5 & 49.0 & 26.2 & 47.4 ($+ 0.9$) \\
  50k & 70.2 & 32.2 & 52.0 & 54.4 & 51.6 & 27.2 & 47.9 ($+ 1.4$)\\
  100k & 70.1 & 33.4 & 52.7 & 55.3 & 52.8 & 27.8 & 48.7 ($+ 2.2$) \\
  500k & \textbf{71.1} & 31.8 & \textbf{53.6} & \textbf{56.5} & \textbf{53.9} & \textbf{28.8} & \textbf{49.3} ($+ 2.8$) \\ \hline
  \end{tabular}
  \caption{The performance on English commonsense reasoning tasks when we use 100B, 200B, and 500B tokens for training.}
  \label{tab:each_data_size_en}
\end{table*}

\begin{table*}[!t]
  \centering
  \footnotesize
  \begin{tabular}{ l | c c c c | c } \hline
  \multicolumn{1}{c|}{\#Vocab} & JSQuAD & JCQA & XWinograd & JAQKET & Avg. \\ \hline \hline
  \multicolumn{6}{c}{10B tokens} \\ \hline \hline
  5k & 1.6 & 37.1 & 51.0 & 0.9 & 22.7 ($\pm 0.0$) \\
  10k & 1.4 & 44.2 & \textbf{53.6} & 0.5 & 24.9 ($+ 2.2$) \\
  50k & 2.7 & 47.9 & 51.0 & 1.6 & 25.8 ($+ 3.1$) \\
  100k & 5.3 & 48.9 & 51.7 & 3.3 & 27.3 ($+ 4.6$) \\
  500k & \textbf{10.1} & \textbf{50.8} & 52.5 & \textbf{4.1} & \textbf{29.4} ($+ 6.7$) \\ \hline \hline
  \multicolumn{6}{c}{50B tokens} \\ \hline \hline
  5k & 36.3 & 49.4 & 53.7 & 3.3 & 35.7 ($\pm 0.0$) \\
  10k & 42.6 & \textbf{59.1} & 56.0 & 7.7 & 41.4 ($+ 5.7$) \\
  50k & 42.8 & 56.8 & 55.7 & 12.2 & 41.9 ($+ 6.2$) \\
  100k & 40.9 & 56.8 & \textbf{56.9} & 17.5 & 43.0 ($+ 7.3$) \\
  500k & \textbf{48.9} & 57.2 & 54.8 & \textbf{17.9} & \textbf{44.7} ($+ 9.0$) \\ \hline \hline
  \multicolumn{6}{c}{100B tokens} \\ \hline \hline
  5k & 45.0 & 55.0 & 56.9 & 5.2 & 40.5 ($\pm 0.0$) \\
  10k & 49.8 & \textbf{60.9} & 56.7 & 12.3 & 44.9 ($+ 4.4$) \\
  50k & 51.4 & 56.0 & 56.8 & 18.4 & 45.7 ($+ 5.2$) \\
  100k & 49.1 & 58.7 & \textbf{58.9} & 20.7 & 46.9 ($+ 6.4$) \\
  500k & \textbf{56.3} & 60.1 & 55.7 & \textbf{26.3} & \textbf{49.6} ($+ 9.1$) \\ \hline \hline
  \multicolumn{6}{c}{200B tokens} \\ \hline \hline
  5k & 50.5 & 58.5 & 56.7 & 7.5 & 43.3 ($\pm 0.0$)\\
  10k & 53.8 & 61.1 & 57.7 & 16.9 & 47.4 ($+ 4.1$)\\
  50k & 55.8 & 54.0 & \textbf{58.7} & 21.4 & 47.5 ($+ 4.2$)\\
  100k & 54.7 & \textbf{64.2} & 58.0 & 27.2 & 51.0 ($+ 7.7$)\\
  500k & \textbf{60.6} & 61.4 & 56.4 & \textbf{32.1} & \textbf{52.6} ($+ 9.3$)\\ \hline \hline
  \multicolumn{6}{c}{500B tokens} \\ \hline \hline  
  5k & 56.4 & 62.2 & 58.4 & 10.4 & 46.9 ($\pm 0.0$) \\
  10k & 59.6 & 62.8 & 58.8 & 20.2 & 50.4 ($+ 3.5$)\\
  50k & 59.3 & \textbf{64.7} & 59.0 & 26.5 & 52.4 ($+ 5.5$)\\
  100k & 60.1 & \textbf{64.7} & \textbf{59.3} & 31.8 & 54.0 ($+ 7.1$)\\
  500k & \textbf{62.6} & 62.9 & 58.8 & \textbf{36.4} & \textbf{55.2} ($+ 8.3$)\\ \hline
  \end{tabular}
  \caption{The performance on Japanese commonsense reasoning tasks when we use 100B, 200B, and 500B tokens for training.}
  \label{tab:each_data_size_ja}
\end{table*}

\end{document}